OpenStudy (pulsified333):

Transplant operations have become routine. One common transplant operation is for kidneys. The most dangerous aspect of the procedure is the possibility that the body may reject the new organ. There are several new drugs available for such circumstances and the earlier the drug is administered, the higher the probability of averting rejection. The New England Journal of Medicine recently reported the development of a new urine test to detect early warning signs that the body is rejecting a transplanted kidney. However, like most other tests, the new test is not perfect. In fact, 20% of negative tests and 6% of positive tests prove to be incorrect. Physicians know that in about 32% of kidney transplants the body tries to reject the organ. If the new test has a positive result (indicating early warning of rejection), what is the probability that the body is attempting to reject the kidney?

OpenStudy (pulsified333):

@satellite73 @dan815

OpenStudy (anonymous):

kind of wordy isn't it? probably a bayes formula question lets see if we can cut through the verbiage

OpenStudy (pulsified333):

okay let me see

OpenStudy (anonymous):

this stuff is usually confusing as hell one way to break through the confusion is to forget about probability and percents, and work with nice large whole numbers to see what is going on

OpenStudy (anonymous):

lets say 10,000 people have the transplant how many try to reject the kidney?

OpenStudy (pulsified333):

3200

OpenStudy (anonymous):

and how many don't ?

OpenStudy (pulsified333):

6800

OpenStudy (anonymous):

20% of negative tests prove to be incorrect. how many is that?

OpenStudy (pulsified333):

of 10000?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

20% of negative tests are incorrect that means out of the 3200 people who DO reject, the test is negative and wrong

OpenStudy (pulsified333):

oh

OpenStudy (pulsified333):

so thats not the answer?

OpenStudy (anonymous):

what is 20% of 3200?

OpenStudy (pulsified333):

640

OpenStudy (pulsified333):

3200-640= 2560?

OpenStudy (anonymous):

those are the people who have a negative nest that is right

OpenStudy (anonymous):

now for the next part

OpenStudy (pulsified333):

so is what I did wrong?

OpenStudy (anonymous):

wait i think i might have done something wrong, the posiive negative business is confusing lets back up a sec

OpenStudy (anonymous):

6800 do not reject 3200 do reject

OpenStudy (anonymous):

out of the the 3200 who do reject, 640 get the wrong answer (negative) and 2560 get the right answer

OpenStudy (pulsified333):

so I have the answer in the work I did earlier

OpenStudy (anonymous):

no not yet, and i still got myself confused, damn we need to work the the 6% now

OpenStudy (anonymous):

ok we are good of the 6800 who do not reject 6% get the wrong answer (positive) so 408 get a positive

OpenStudy (anonymous):

now we can finish (thank god)

OpenStudy (anonymous):

If the new test has a positive result (indicating early warning of rejection), what is the probability that the body is attempting to reject the kidney? it is the number of people who do reject and test positive, divided by all that test positive

OpenStudy (anonymous):

the number that test positive (if i did it correctly) is 2560+408

OpenStudy (pulsified333):

3608

OpenStudy (anonymous):

and the number of people who test positive and have it is 2560

OpenStudy (pulsified333):

okay

OpenStudy (anonymous):

now we can try this using bayes formula and not numbers if you like just working with the percents as decimals

OpenStudy (pulsified333):

Pr(A|B) = (Pr(B|A)Pr(A))/Pr(B)

OpenStudy (anonymous):

we need the percent of people that you test positive and have it, which is \[.32\times .80\]

OpenStudy (anonymous):

that is your numerator

OpenStudy (anonymous):

and you need percent who test positive \[.32\times .80+.68\times .06\]

OpenStudy (anonymous):

should be the same answer

OpenStudy (pulsified333):

as?

OpenStudy (pulsified333):

Which one?

OpenStudy (anonymous):

i made a typo in the first one http://www.wolframalpha.com/input/?i=%28.32*.80%29%2F%28.32*.80%2B.06*.68%29

OpenStudy (anonymous):

they are the same

OpenStudy (pulsified333):

I see

OpenStudy (pulsified333):

:D it worked thanks.

OpenStudy (pulsified333):

I have to finish as fast as I can so that I can study for my midterm tomorrow